mary is making quilts for her family. Each quilt will have 12 triangles. The triangles have a base 4in and a height of 8 in. She will be making 9 quilts. how much material will she need to make all the triangles?
Answer by jim_thompson5910(35256)


Area of One Triangle = 0.5*b*h
Area of One Triangle = 0.5*4*8
Area of One Triangle = 2*8
Area of One Triangle = 16 square inches
------------------------------------
Area of one quilt = Area of 12 triangles
Area of one quilt = 12*(area of one triangle)
Area of one quilt = 12*(16)
Area of one quilt = 192 square inches
------------------------------------
Area of 9 quilts = 9*(area of one quilt)
Area of 9 quilts = 9*(192)
Area of 9 quilts = 1728
------------------------------------
------------------------------------
Final Answer: Mary will need 1728 square inches of material.